# NeRF-CA: Dynamic Reconstruction of X-ray Coronary Angiography with Extremely Sparse-views

### v1

@@ V Experiments
In this section, we present experiments to evaluate and validate NeRF-CA. We present the 4D phantom datasets that allow us to evaluate quantitatively. Our method has multiple components and parameters that we evaluate through extensive ablation studies. Furthermore, we compare our method to existing state-of-the-art to show the strengths and weaknesses of our approach.
@@ V-A Datasets
For the evaluation of NeRF-CA, we rely on parameterized 4D synthetic phantoms that have been used before in a similar evaluation context [45]. These phantoms allow us to generate realistic ground-truth X-ray sequences from the anatomical cardiac vessel structure from any viewpoint at any cardiac motion phase. Utilizing these phantoms, we can effectively focus our evaluation on the aspects NeRF-CA addresses, i.e., vessel sparsity, background occlusion, and the modeling of synchronized cardiac motion. We note that these phantoms do not exhibit other complex characteristics that are present in real clinical CA data, such as blood vessel contrast inhomogeneity, lack of cardiac synchronization among views, and respiratory motion, which are not addressed by NeRF-CA. Clinical CA data exhibiting all the complex characteristics, as described in Section II-A, would limit the value of the evaluation through the confounding factors influencing the results. Furthermore, the flexibility for ablation studies is rather limited on real clinical CA data, for example, we would like to utilize more CA sequences than the maximum of $4$ acquired per patient in the clinical setting. Moreover, no 3D or 4D ground-truth volumes are available in clinical CA data that from which we can generate or evaluate 2D CA sequences. Therefore, we use the 4D XCAT phantom [13] and a 4D coronary computed tomography angiography (CCTA) scan, called MAGIX [14]. Both datasets depict the left coronary artery (LCA), which is more complex in geometry than the right coronary artery and, therefore, more clinically interesting for 3D reconstruction applications. For both datasets, we simulate homogeneous contrast injection using higher attenuation values in segmented areas, similar to previous work [6]. Figure 1 shows examples of one angiogram sequence for these 4D datasets. In these datasets, one angiogram consists of $T=10$ frames, each corresponding to a discrete cardiac phase $i$, assuming respiratory motion to be minimal. As can be seen from Figure 1, the 4D XCAT dataset [13] includes a detailed LCA structure, fitting for evaluating the effect of vessel sparsity on our reconstruction quality. Moreover, the 4D XCAT phantom allows for the simulation of cardiac motion according to the ECG-based cardiac phase. We sample the XCAT phantom with an isotropic resolution of 0.5 mm3. The 4D CCTA dataset MAGIX, obtained from the OsiriX platform [14], consists of 10 consecutive scans that represent different cardiac phases across the cardiac cycle. As contrast is injected throughout the body during CCTA scans, each scan also includes varying contrast in the background, which allows us to evaluate the effect of the dynamic background on our reconstructions. Due to the CCTA’s lower resolution of $0.4\times 0.4\times 2$ mm3, the obtained LCA structure is coarse. To generate projections from the 4D volumes, we utilize the tomographic toolbox TIGRE [46]. We adjust the geometry settings to simulate the C-arm setting of CA data, where the focus is on the main branches while deeper branches may be out of view.
@@
As discussed in Section II-A, the viewing angles in CA are literature-based optimal angles to minimize overlap and foreshortening [17]. The LCA is typically visualized with 8 distinctive angiographic views [47, 48]. Figure 4 shows these views in Euler angles ($\theta$, $\phi$) distributed along a flattened sphere. The image examples are from the XCAT dataset. For 4D reconstruction, selecting viewing angles in a sparse setting is crucial. We, therefore, base our training and validation sets on these optimal views. We will train our method using $4$ to $70$ projections. For $4$ projections, we choose $4$ well-distributed optimal angles, indicated by the yellow circles in the figure and images (a-d). The remaining views, indicated by pink circles and images (e-h), are used to validate all cases. For more projections, we uniformly sample training views within a limited range of $60^{\circ}$, indicated by the light yellow area. This area covers the range of most optimal viewing angles while avoiding views too close to the validation set.
@@ V-B Evaluation set-up
We perform both qualitative and quantitative evaluations in our work. The qualitative approach is suitable to overcome known problems in assessing medical imaging with pure quantitative metrics [49]. For our application, where topological accuracy is most relevant, this is especially relevant [6]. We evaluate quantitatively by computing the re-projection error for 2D novel view synthesis, fitting the roadmap application where the 4D reconstruction would be re-projected to the C-arm angle of interest.
The ultimate goal of our work is to evaluate the reconstruction of the coronary artery structure rather than direct volume-to-volume comparison. Specifically, we are interested in the topology of the structure, for which, to our knowledge, no appropriate quantitative measures have been proposed [6]. Therefore, we mostly rely on qualitative analysis for the structure itself. Besides, we also perform binary comparisons by computing Dice, a well-known metric for blood vessel segmentation tasks [50]. We only consider the positive or blood vessel structure class when computing Dice. This approach compensates for the imbalance of these classes due to the sparsity of the structure in the image. To compute a binary Dice value, we generate novel views by applying maximum intensity projection (MIP) from the dynamic attenuation values $\sigma_{d}$, assuming that the blood vessel structure has the highest attenuation values and is purely dynamic. This means that we strictly consider the dynamic coronary artery component for our Dice computation. We obtain the final binarized images from thresholding the MIP views based on the 4D ground-truth values of the coronary artery structure. When comparing to static models, we compute the MIP from the static $\sigma$ attenuation values. We also evaluate the overall 2D CA image prediction with common image quality metrics peak-signal-to-noise ratio (PSNR) and structural similarity (SSIM). Besides quality metrics, we also report the computation time.
@@ V-C Implementation Details
The models [^1] https://github.com/kirstenmaas/NeRF-CA were implemented in Pytorch. The models were trained on an RTX A5000 GPU. We utilize the same hyperparameters for both datasets. The models have an MLP architecture of $4$ hidden layers with $128$ neurons with ReLU activations, commonly used in NeRF applications [5]. We utilize a Softplus activation to obtain a positive attenuation value output. We train our models with an Adam optimizer with a linearly decaying learning rate from $1\times 10^{-3}$ to $1\times 10^{-5}$ in $150000$ steps. The number of training iterations is $200000$ with a ray batch size of $1024$. We train with images of size $200\times 200$ pixels and sample $500$ points along each ray.
The hyperparameters for our method components were chosen based on a grid search. Although many different hyperparameters are involved, we show they can be consistently used across different datasets. We impose a maximum frequency band for the windowed positional encoding, as shown in Equation 4, of $L=12$ over $N=150000$ iterations. We start the coarse-to-fine transition at a frequency band of $1$, therefore activating the first frequency bands $(\sin(\pi\boldsymbol{x}),\cos(\pi\boldsymbol{x}))$ from the start of training. This has shown to be more advantageous for separating foreground and background. A fraction of $V=0.5$ of the ray batch size is sampled for the purpose of weighted pixel sampling. The static vs. dynamic factorization loss weight $\lambda_{b}$ increases from $1\times 10^{-12}$ to $1\times 10^{-10}$, with a delay of $40000$ iterations. The dynamic entropy loss weight $\lambda_{e}$ linearly increases from $1\times 10^{-12}$ to $1\times 10^{-10}$ without a delay, where only rays are considered with a minimum accumulated density of $1\times 10^{-4}$. The dynamic occlusion loss weight $\lambda_{o}$ increases from $1\times 10^{-8}$ to $1\times 10^{-5}$ linearly, delayed by $40000$ iterations with a distance threshold of $D=0.2$.
@@ V-D Comparison to Existing Methods
In this section, we compare our method to state-of-the-art 3D reconstruction methods. As discussed in Section III-A, coronary X-ray angiography techniques require manual interactions or large training datasets. As our method does not require these settings, we compare it to state-of-the-art NeRF X-ray reconstruction methods, which obtain static reconstructions from one dataset directly. We particularly focus on state-of-the-art methods for sparse-view reconstruction, according to the main goal of our work. We compare our method to SAX-NeRF [36] and NAF [37], as well as the original NeRF method as a baseline [5]. Lastly, we include our method with only the sparse-view reconstruction components, referred to as sparse method $\mathcal{L}_{s}$, as a representative of a state-of-the-art sparse-view natural scene NeRF-based reconstruction method [8]. This sparse method is represented by the loss $\mathcal{L}_{s}$, where $\mathcal{L}_{s}(\boldsymbol{r},\boldsymbol{\tau}_{i})=\mathcal{L}_{p}(\boldsymbol{r},\boldsymbol{\tau}_{i})$. From our proposed components, only the photometric loss $\mathcal{L}_{p}$ (see Equation 1) and the windowed positional encoding $\gamma$ (see Equation 4) is applied. We utilize one static MLP trained individually for every single timestep. We note that, unlike the natural scene method [8], we do not apply the dynamic occlusion loss $\mathcal{L}_{o}$ (see Equation 5). This is because the assumption of edge emptiness, as described in Section IV-D, does not apply to our static non-empty scene.
@@ V-E Ablation Studies
We evaluate the multiple components and parameters of our work through ablation studies. Firstly, we examine the dynamic and sparse-view reconstruction components individually, as introduced in Figure 2 (a). Next, we analyze the consistency in quality for the dynamic reconstruction across the cardiac phases. Lastly, we evaluate the effect of weighted pixel sampling and the dynamic entropy and occlusion loss components.
Dynamic and Sparse-view Reconstruction Components    In Section IV and Figure 2 (a), we illustrate how NeRF-CA effectively combines dynamic and sparse-view reconstruction components. In this section, we discuss the outcomes of our full method $\mathcal{L}_{f}$ compared to these separate components, denoted as our sparse method and our dynamic method, respectively. The definition of our sparse method with loss $\mathcal{L}_{s}$ is given in Section V-D. We define our dynamic method, as shown in Figure 2 (a) in green, with loss $\mathcal{L}_{d}$ as
@@
where we also apply the windowed positional encoding $\gamma$ (see Equation 4. We exclude the dynamic occlusion loss $\mathcal{L}_{o}$ to verify its role in sparse-view reconstruction. Figure 6 shows the quantitative results for the XCAT and MAGIX dataset, where (a-c) show the Dice, PSNR, and SSIM scores for the XCAT dataset and (d-f) show these scores for the MAGIX dataset. We display these scores across various projection counts for our full $\mathcal{L}_{f}$, sparse $\mathcal{L}_{s}$, and dynamic $\mathcal{L}_{d}$ methods. The low-opacity lines represent the scores of the four validation views averaged over all timesteps, while the bold line shows the mean score across these views.
@@
First, we evaluate the reconstruction of only the coronary artery structure using Dice scores, as it addresses the main goal of our work. Our full method $\mathcal{L}_{f}$ achieves high Dice scores in the extremely sparse-view settings of 4 and 9 projections. These scores are consistent across the four validation views, as shown by the low-opacity lines. For the XCAT dataset, our full method $\mathcal{L}_{f}$ significantly outperforms our sparse method $\mathcal{L}_{s}$ and slightly outperforms our dynamic method $\mathcal{L}_{d}$. The scores across the MAGIX dataset are more similar, showing that all three of our methods perform well on this dataset, although the sparse method $\mathcal{L}_{s}$ slightly outperforms our other methods. Figure 7 shows the qualitative examples for the $4$ training projections setting. We show the static output, dynamic output, composite output, and ground truth images. From these qualitative examples, we can observe that our full method $\mathcal{L}_{f}$ achieves the best separation of the dynamic coronary artery from the static background, removing the occluding background noise from the novel views. For the XCAT dataset, the reconstructed structure is topologically accurate, with the primary branches clearly defined, which fits our clinical applications. For the MAGIX dataset, our models achieve a high-fidelity structure without loss of quality, likely due to the easier coarser nature of the structure. We also report scores beyond the extremely sparse-view setting. For many projections, our full $\mathcal{L}_{f}$, dynamic $\mathcal{L}_{d}$, and sparse $\mathcal{L}_{s}$ methods all perform well in reconstructing the coronary artery structure. However, our full $\mathcal{L}_{f}$ and dynamic $\mathcal{L}_{d}$ methods lose quality with more training projections, whereas our sparse method $\mathcal{L}_{s}$ does not. Our full method $\mathcal{L}_{f}$ and dynamic $\mathcal{L}_{d}$ method impose regularization constraints on the separation of foreground and background, which may explain the score loss with more projections. Although our work focuses on extremely sparse-view settings, the sparse method $\mathcal{L}_{s}$ might be more suitable for reconstructions with more projections.
@@
We also report the overall CA image comparison scores in PSNR and SSIM in Figure 6 and the overall running time. The PSNR scores indicate that although our full model is more susceptible to background noise, the SSIM scores indicate that the overall performance is still accurate, even in the extremely sparse setting of $4$ and $9$ projections for both datasets. The sparse method $\mathcal{L}_{s}$ is more accurate in background reconstruction when presented with more projections, highlighting its potential outside the extremely sparse-view setting. We do observe variation in scores among the validation views, which is likely due to background occlusion, which can be observed in the qualitative examples. However, separating the structure from the background mitigates this issue. Therefore, our full method $\mathcal{L}_{f}$ remains effective, whereas our sparse $\mathcal{L}_{s}$ and dynamic $\mathcal{L}_{d}$ method may be less applicable due to background occlusion. Since the Dice score best represents our expected results and addresses this separation, we will only report this score for the remaining cases. The running time of our full method $\mathcal{L}_{f}$, dynamic method $\mathcal{L}_{d}$, and sparse method $\mathcal{L}_{f}$ are $5.5$, $5.5$ and $3$ hours, respectively.
@@
Figure 8 shows how our method compares to the ground truth for our two datasets, (a) XCAT and (b) MAGIX, in a setting of $4$ angiogram sequences. We display the scores for the four validation views as low opacity lines, and the mean score computed from these views as a bold line. For the (a) XCAT dataset, the line graphs indicate that the scores of our full $\mathcal{L}_{f}$ method are consistent across the cardiac phases, which can be derived from the constant slope of the line. The qualitative examples, displaying the dynamic output and MIP of the ground truth, show that the slightly lower quality of the full method $\mathcal{L}_{f}$ is due to the missing sparse branches. These sparse branches are less important in our application, where topology is most relevant. Therefore, our full method $\mathcal{L}_{f}$ maintains quality across the dynamic cardiac phases. For the (b) MAGIX dataset, we observe more variation in Dice score across timesteps and among the different validation views. We can observe that this variation is due to background noise appearing in the dynamic scene. The MAGIX dataset consists of dynamic contrast flow across the cardiac phases, resulting in these parts being modeled as dynamic by our model. However, the background noise is primarily outside of the coronary artery structure, not affecting the reconstruction quality of the structure itself. All in all, our model consistently maintains quality for reconstructing the coronary artery structure, fitting our clinical application.
@@
Figure 8: Dice scores for our full method $\mathcal{L}_{f}$ over the cardiac phases for the (a) XCAT and (b) MAGIX datasets. The model is trained with $4$ angiogram sequences. The low-opacity lines represent the scores of the four validation views, whereas the bold line represents the mean score across these views.
@@
Weighted Pixel Sampling   We incorporate weighted pixel sampling in our method to enhance the reconstruction of blood vessel structures. This approach targets sampling in the heuristically determined regions of the blood vessels, enhancing the reconstruction of these sparse structures, as described in Section IV-D. Figure 9 (a) displays the Dice scores for our full method $\mathcal{L}_{f}$ and our method without weighted pixel sampling (w/o wps) over the number of training projections for the XCAT dataset. The results indicate that weighted pixel sampling generally improves the score across all projection counts, significantly boosting performance in the extremely sparse setting of $4$ projections from a nearly zero score to a high Dice score with accurate structure reconstruction. For the MAGIX dataset, we also observe a slight overall score increase with weighted pixel sampling, as shown in the supplementary material.
@@
Dynamic Entropy and Occlusion Loss    We introduce the dynamic entropy $\mathcal{L}_{e}$ and dynamic occlusion $\mathcal{L}_{o}$ losses to accurately decouple the scene in a static background and dynamic coronary artery, as described in Section IV-D and Section IV-E, respectively. Since both losses affect the dynamic model, we analyze their combined effectiveness, as they are likely interrelated. Figure 9 (b) shows the Dice scores for the XCAT dataset across the dynamic entropy $\lambda_{e}$ and dynamic occlusion $\lambda_{o}$ weights for the 4 training projection settings. The optimal weights for our model are indicated by the black square with an entropy weight $\lambda_{e}$ of $e^{-10}$ and the occlusion weight $\lambda_{o}$ of $e^{-5}$, leading to a Dice score of $0.78$. We observe that the most optimal Dice score is obtained by combining both losses, demonstrating their effectiveness. For the MAGIX dataset, we observe similar patterns for the same hyperparameters, which are shown in the supplementary material.

### v2

@@ V Experiments
In this section, we present experiments to evaluate and validate NeRF-CA. We present the 4D phantom datasets that allow us to evaluate quantitatively. Our method has multiple components and parameters that we evaluate through extensive ablation studies. Furthermore, we compare our method to existing state-of-the-art methods to show the strengths and weaknesses of our approach.
@@ V-A Datasets
For the evaluation, we use parameterized 4D synthetic phantoms previously used in a similar evaluation context [43]. These phantoms allow us to generate realistic ground-truth X-ray sequences from the anatomical cardiac vessel structure from any viewpoint at any cardiac phase. Utilizing these phantoms, we can effectively focus our evaluation on the aspects NeRF-CA specifically addresses, i.e., sparse-views and cardiac motion. Meanwhile, these phantoms also exhibit varying aspects of vessel sparsity and poor vessel visibility, allowing us to validate our model under these conditions. These phantoms intentionally exclude complex characteristics of real clinical CA data, such as blood vessel contrast inhomogeneity, lack of cardiac synchronization among views, and respiratory motion, to ensure the evaluation remains focused on the aspects NeRF-CA is designed to address. Using clinical data with these factors, as described in Section II-A, would introduce confounding variables, limiting the evaluation’s value. Additionally, real clinical CA data offers limited flexibility for ablation studies; for example, the typical clinical setting provides a maximum of 4 views per patient. Moreover, clinical CA data lacks 3D or 4D ground-truth volumes needed for generating or evaluating 2D CA sequences.
@@
We use the 4D XCAT phantom [16] and a 4D coronary computed tomography angiography (CCTA) scan, called MAGIX [17]. Both datasets depict the left coronary artery (LCA), which is more complex in geometry than the right coronary artery and, therefore, more clinically interesting for 3D reconstruction applications. For both datasets, we simulate homogeneous contrast injection using higher attenuation values in segmented areas, similar to previous work [7]. Figure 1 shows examples of one angiogram sequence for these 4D datasets. In these datasets, one angiogram consists of $T=10$ frames, each corresponding to a discrete cardiac phase $i$, assuming minimal respiratory motion. As can be seen from Figure 1, the 4D XCAT dataset [16] includes a detailed LCA structure, fitting for evaluating the effect of vessel sparsity on our reconstruction quality. Moreover, the 4D XCAT phantom allows for the simulation of cardiac motion according to the ECG-based cardiac phase. We sample the XCAT phantom with an isotropic resolution of 0.5 mm3. The 4D CCTA dataset MAGIX, obtained from the OsiriX platform [17], consists of 10 consecutive scans that represent different cardiac phases across the cardiac cycle. As contrast is injected throughout the body during CCTA scans, each scan also includes varying contrast in the background, which allows us to evaluate the effect of the dynamic background on our reconstructions. Due to the CCTA’s lower resolution of $0.4\times 0.4\times 2$ mm3, the obtained LCA structure is coarse. To generate projections from the 4D volumes, we utilize the tomographic toolbox TIGRE [44]. We adjust the geometry settings to simulate the C-arm setting of CA data, where the focus is on the main branches while deeper branches may be out of view.
@@
In clinical practice, the LCA is typically visualized with 8 literature-based distinctive angiographic views [45, 46] to minimize overlap and foreshortening [47]. Figure 3 shows these views in Euler angles ($\theta$, $\phi$) distributed along a flattened sphere. The image examples are from the XCAT dataset. For 4D reconstruction, selecting viewing angles in a sparse setting is crucial. We, therefore, base our training and validation sets on these clinical optimal views. We will train our method using $4$ to $70$ projections. For $4$ projections, we choose $4$ well-distributed optimal angles, indicated by the yellow circles in the figure and images (a-d). The remaining views, indicated by pink circles and images (e-h), are used to validate all cases. For more projections, we uniformly sample training views within a limited range of $60^{\circ}$, indicated by the light yellow area. This area covers the range of most optimal viewing angles while avoiding views too close to the validation set.
@@ V-B Evaluation set-up
We perform both qualitative and quantitative evaluations in our work. The qualitative approach is suitable to overcome known problems in assessing medical imaging with pure quantitative metrics [48]. For our application, where topological accuracy is most relevant, this is especially valuable [7]. We evaluate quantitatively by computing the re-projection error for 2D novel view synthesis, fitting the roadmap application where the 4D reconstruction would be re-projected to the C-arm angle of interest, as discussed in Section II-A.
The ultimate goal of our work is to evaluate the reconstruction of the coronary artery structure rather than direct volume-to-volume comparison. Specifically, we are interested in the topology of the structure, for which, to our knowledge, no appropriate quantitative measures have been proposed [7]. Therefore, we mostly rely on qualitative analysis for the structure itself. Besides, we also perform binary comparisons by computing Dice, a well-known metric for blood vessel segmentation tasks [49]. We only consider the positive or blood vessel structure class when computing Dice. This approach compensates for the imbalance of these classes due to the sparsity of the structure in the image. To compute a binary Dice value, we generate novel views by applying maximum intensity projection (MIP) from the dynamic attenuation values $\sigma_{d}$, assuming that the blood vessel structure has the highest attenuation values and is purely dynamic. This means that we strictly consider the dynamic coronary artery component for our Dice computation. We obtain the final binarized images from thresholding the MIP views based on the 4D ground-truth values of the coronary artery structure. When comparing to static models, we compute the MIP from the composite $\sigma$ attenuation values. We also evaluate the overall 2D CA image prediction with common image quality metrics peak-signal-to-noise ratio (PSNR) and structural similarity (SSIM). Besides quality metrics, we also report the computation time.
@@ V-C Implementation Details
The models were implemented in Pytorch and trained on an RTX A5000 GPU. We utilize the same hyperparameters for both datasets. The models have an MLP architecture of $4$ hidden layers with $128$ neurons with ReLU activations, commonly used in NeRF applications [6]. We utilize a Softplus activation to obtain a positive attenuation value output. We train our models with an Adam optimizer with a linearly decaying learning rate from $1\times 10^{-3}$ to $1\times 10^{-5}$ in $150000$ steps. The number of training iterations is $200000$ with a ray batch size of $1024$. We train with images of size $200\times 200$ pixels and sample $500$ points along each ray.
The hyperparameters for our method components were chosen based on a grid search. Although many different hyperparameters are involved, we show they can be consistently used across different datasets, highlighting the robustness of our method. We impose a maximum frequency band for the windowed positional encoding, as shown in Equation 5, of $L=12$ over $N=150000$ iterations. We start the coarse-to-fine transition at a frequency band of $1$, therefore activating the first frequency bands $(\sin(\pi\boldsymbol{x}),\cos(\pi\boldsymbol{x}))$ from the start of training. This has shown to be more advantageous for separating foreground and background. A fraction of $V=0.5$ of the ray batch size is sampled for the purpose of weighted pixel sampling. The static vs. dynamic factorization loss weight $\lambda_{b}$ increases from $1\times 10^{-12}$ to $1\times 10^{-10}$, with a delay of $40000$ iterations. The dynamic entropy loss weight $\lambda_{e}$ linearly increases from $1\times 10^{-12}$ to $1\times 10^{-10}$ without a delay, where only rays are considered with a minimum accumulated density of $1\times 10^{-4}$. The dynamic occlusion loss weight $\lambda_{o}$ increases from $1\times 10^{-8}$ to $1\times 10^{-5}$ linearly, delayed by $40000$ iterations with a distance threshold of $D=0.2$.
@@ V-D Comparison to Existing Methods
In this section, we compare our method to state-of-the-art 3D reconstruction methods. As discussed in Section III-A, coronary X-ray angiography techniques require manual interactions or large training datasets. As our method does not require these settings, we compare it to state-of-the-art NeRF X-ray reconstruction methods, which obtain static reconstructions from one dataset directly. We particularly focus on state-of-the-art methods for sparse-view reconstruction, according to the main goal of our work. We compare our method to SAX-NeRF [10] and NAF [36], as well as the original NeRF method as a baseline [6]. Lastly, we include our method with only the sparse-view reconstruction components, referred to as sparse method $\mathcal{L}_{s}$, as a representative of a state-of-the-art sparse-view natural scene NeRF-based reconstruction method [9]. This sparse method is represented by a part of our full loss $\mathcal{L}_{f}$ (see Equation 1) as loss $\mathcal{L}_{s}$, where $\mathcal{L}_{s}(\boldsymbol{r},\boldsymbol{\tau}_{i})=\mathcal{L}_{p}(\boldsymbol{r},\boldsymbol{\tau}_{i})$. From our proposed components, only the photometric loss $\mathcal{L}_{p}$ (see Equation 2) and the windowed positional encoding $\gamma$ (see Equation 5) is applied. We utilize one static MLP trained individually for every single timestep. We note that, unlike the natural scene method [9], we do not apply the dynamic occlusion loss $\mathcal{L}_{o}$ (see Equation 6). This is because the assumption of edge emptiness, as described in Section IV-D, does not apply to our static non-empty scene.
@@ V-E Ablation Studies
We evaluate the multiple components and parameters of our work through ablation studies. Firstly, we examine the dynamic and sparse-view reconstruction components individually, as introduced in Figure 2 (a). Next, we analyze the consistency in quality for the dynamic reconstruction across the cardiac phases. Lastly, we evaluate the effect of weighted pixel sampling and the dynamic entropy and occlusion loss components. Furthermore, we validate our choice for self-supervision over a segmentation-supervised approach, which can be found in the supplementary material. Specifically, we demonstrate that the self-supervised approach achieves higher accuracy, as it does not depend on the automatic error-prone segmentation methods that produce inconsistencies in supervision.
Dynamic and Sparse-view Reconstruction Components    In Section IV and Figure 2 (a), we illustrate how NeRF-CA effectively combines dynamic and sparse-view reconstruction components. In this section, we discuss the outcomes of our full method $\mathcal{L}_{f}$ compared to these separate components, denoted as our sparse method and our dynamic method, respectively. The definition of our sparse method with loss $\mathcal{L}_{s}$ is given in Section V-D. We define our dynamic method, as shown in Figure 2 (a) in green, as a part of our full loss $\mathcal{L}_{f}$ (see Equation 1) as loss $\mathcal{L}_{d}$
@@
where we also apply the windowed positional encoding $\gamma$ (see Equation 5). We exclude the dynamic occlusion loss $\mathcal{L}_{o}$ (see Equation 1) to verify its role in sparse-view reconstruction. Figure 5 shows the quantitative results for the XCAT and MAGIX dataset, where (a-c) show the Dice, PSNR, and SSIM scores for the XCAT dataset and (d-f) show these scores for the MAGIX dataset. We display these scores across various projection counts for our full $\mathcal{L}_{f}$ (in orange), sparse $\mathcal{L}_{s}$ (in blue), and dynamic $\mathcal{L}_{d}$ (in green) methods. The low-opacity lines represent the scores of the four validation views averaged over all timesteps, while the bold line shows the mean score across these views.
@@
First, we evaluate the reconstruction of only the coronary artery structure using Dice scores, as it addresses the main goal of our work. Our full method $\mathcal{L}_{f}$ (orange) achieves high Dice scores in the extremely sparse-view settings of 4 and 9 projections. These scores are consistent across the four validation views, as shown by the low-opacity lines. For the XCAT dataset, our full method $\mathcal{L}_{f}$ significantly outperforms our sparse method $\mathcal{L}_{s}$ (blue) and slightly outperforms our dynamic method $\mathcal{L}_{d}$ (green), as can be seen from Figure 5 (a). The scores across the MAGIX dataset are more similar, as seen in Figure 5 (d), showing that all three of our methods perform well on this dataset, although the sparse method $\mathcal{L}_{s}$ slightly outperforms our other methods. Figure 6 shows the qualitative examples for the $4$ training projections setting. We show the static output, dynamic output, composite output, and ground truth images. By comparing the different methods across the rows, we can observe that our full method $\mathcal{L}_{f}$ (bottom row, orange) achieves the best separation of the dynamic coronary artery from the static background, removing the occluding background from the novel views. For the XCAT dataset (Figure 6 (a)), the reconstructed structure is topologically accurate, with the primary branches clearly defined, which fits our clinical applications. For the MAGIX dataset (Figure 6 (b)), our models achieve a high-fidelity structure without loss of quality, likely due to the easier coarser nature of the structure. We also report scores beyond the extremely sparse-view setting, as seen in Figure 5(a,d). For many projections, our full $\mathcal{L}_{f}$, dynamic $\mathcal{L}_{d}$, and sparse $\mathcal{L}_{s}$ methods all perform well in reconstructing the coronary artery structure. However, our full $\mathcal{L}_{f}$ and dynamic $\mathcal{L}_{d}$ methods lose quality with more training projections, whereas our sparse method $\mathcal{L}_{s}$ does not as can be seen from the flat line. Our full method $\mathcal{L}_{f}$ and dynamic $\mathcal{L}_{d}$ method impose regularization constraints on the separation of foreground and background, which may explain the score loss with more projections. Although our work focuses on extremely sparse-view settings, the sparse method $\mathcal{L}_{s}$ might be more suitable for reconstructions with more projections.
@@
We also report the overall CA image comparison scores in PSNR and SSIM in Figure 5 and the overall running time. The drop in PSNR scores (Figure 6 (b, e)) indicates that our full method is susceptible to background noise. Meanwhile, the SSIM scores (Figure 6 (c, f)) indicate that the overall performance is still accurate, even in the extremely sparse setting of $4$ and $9$ projections for both datasets. The sparse method $\mathcal{L}_{s}$ (blue) is more accurate in background reconstruction when presented with more projections, highlighting its potential outside the extremely sparse-view setting. We do observe variation in scores among the validation views as can be seen from the low-opacity lines in Figure 5 (b,c,e,f), which is likely due to poor vessel visibility, which can be observed in the qualitative examples in Figure 6. However, separating the structure from the background mitigates this issue. Therefore, our full method $\mathcal{L}_{f}$ remains effective, whereas our sparse $\mathcal{L}_{s}$ and dynamic $\mathcal{L}_{d}$ methods may be less applicable due to poor vessel visibility. Since the Dice score best reflects our application, which is reconstructing the blood vessel structure, we will only report this score for the remaining cases. The running time of our full method $\mathcal{L}_{f}$, dynamic method $\mathcal{L}_{d}$, and sparse method $\mathcal{L}_{f}$ are $5.5$, $5.5$ and $3$ hours, respectively.
@@
Figure 7 shows how our method compares to the ground truth for our two datasets, (a) XCAT and (b) MAGIX, in a setting of $4$ angiogram views. We display the scores for the four validation views as low opacity lines, and the mean score computed from these views as a bold line. For the (a) XCAT dataset, the line graphs indicate that the scores of our full $\mathcal{L}_{f}$ method are consistent across the cardiac phases, which can be derived from the constant slope of the line. The qualitative examples, displaying the dynamic output and MIP of the ground truth, show that the slightly lower quality of the full method $\mathcal{L}_{f}$ is due to the missing sparse branches. These sparse branches are less important in our application, where topology is most relevant. Therefore, our full method $\mathcal{L}_{f}$ maintains quality across the dynamic cardiac phases. For the (b) MAGIX dataset, we observe more variation in Dice score across timesteps and among the different validation views. We can observe that this variation is due to background noise appearing in the dynamic scene. The MAGIX dataset consists of dynamic contrast flow across the cardiac phases, resulting in these parts being modeled as dynamic by our model. However, the background noise is primarily outside of the coronary artery structure, not affecting the reconstruction quality of the structure itself. All in all, our model consistently maintains quality for reconstructing the coronary artery structure, fitting our clinical application.
@@
Figure 7: Dice scores for our full method $\mathcal{L}_{f}$ over the cardiac phases for the (a) XCAT and (b) MAGIX datasets. The model is trained with $4$ angiogram views. The low-opacity lines represent the scores of the four validation views, whereas the bold line represents the mean score across these views.
@@
Weighted Pixel Sampling   We incorporate weighted pixel sampling in our method to enhance the reconstruction of blood vessel structures. This approach targets sampling in the heuristically determined regions of the blood vessels, enhancing the reconstruction of these sparse structures, as described in Section IV-D. Figure 8 (a) displays the Dice scores for our full method $\mathcal{L}_{f}$ and our method without weighted pixel sampling (w/o wps) over the number of training projections for the XCAT dataset. The results indicate that weighted pixel sampling (in orange in Figure 8 (a)) generally improves the score across all projection counts, significantly boosting performance in the extremely sparse setting of $4$ projections from a nearly zero score to a high Dice score with accurate structure reconstruction. For the MAGIX dataset, we also observe a slight overall score increase with weighted pixel sampling, as shown in the supplementary material.
Dynamic Entropy and Occlusion Loss    In this section, we evaluate the the dynamic entropy $\mathcal{L}_{e}$ and dynamic occlusion $\mathcal{L}_{o}$ losses, which are described in Section IV-D and Section IV-E, respectively. Since both losses affect the dynamic model, we analyze their combined effectiveness, as they are likely interrelated. Figure 8 (b) shows the Dice scores for the XCAT dataset across the dynamic entropy $\lambda_{e}$ and dynamic occlusion $\lambda_{o}$ weights for the 4 training projection settings. The optimal weights for our model are indicated by the black square with an entropy weight $\lambda_{e}$ of $e^{-10}$ and the occlusion weight $\lambda_{o}$ of $e^{-5}$, leading to a Dice score of $0.78$. We observe that the most optimal Dice score is obtained by combining both losses, demonstrating that both are essential for optimal results. For the MAGIX dataset, we observe similar patterns for the same hyperparameters, as shown in the supplementary material.